OpenStudy (anonymous):

Solve this out for me pleasee –2(x – 3) ≤ 5 – (x + 3)

OpenStudy (anonymous):

-2(x-3)=5-(x+3) -2x+6=-x+2 -2x+x=2-6 -x=-4 x=4

OpenStudy (anonymous):

Thanks (: